A body that started from rest moves with constant acceleration in a straight line. After travelling a distance d the speed of the car is v. What is the distance travelled when the speed of the car was v/2? (Ans: d/4)


1
Expert's answer
2022-04-26T14:56:26-0400

"v^2\sim d,"

"(\frac v{\frac v2})^2=\frac dx,"

"x=\frac d4 ."